# Giovanni Compiani Ilya Morozov Stephan Seiler

### v3

@@ 4.3 Relevance of Different Data Types
Lastly, we analyze which types of unstructured data yield the largest fit improvements. It would be misleading to report only the best-performing model in each category, as multiple text or image models may achieve similar $AIC$ improvements over plain logit. We therefore follow the statistical literature on model selection and report Akaike weights (Burnham and Anderson, 2004). Formally, for each data type $d$ (images, titles, descriptions, or reviews) we compute the Akaike weight $w_{d}$ as
$w_{d}=\frac{\sumop\displaylimits_{\{r:d(r)=d\}}\exp(-{}_{r}/2)}{\sumop\displaylimits_{i}\exp(-{}_{i}/2)}$ (3)
where $i$ indexes estimated specifications, and $_{i}=AIC_{i}-AIC_{min}$ where $AIC_{min}$ is the lowest $AIC$ across all estimated specifications in that product category. The numerator in (3) sums over estimated specifications using data type $d$, while the denominator sums over all specifications estimated during model selection. We interpret $w_{d}$ heuristically as the posterior probability that the mixed logit model based on data type $d$ is the best model given the data. [^26] More precisely, in large samples, $w_{d}$ reflects the probability that this class of models is, in fact, the best model for the data in the sense of Kullback-Leibler information (Burnham and Anderson, 2004, p.272).
Figure 8: Akaike Weights of different types of unstructured data across 40 product categories. These weights reflect the relative importance of data types for predicting substitution patterns in the data. We compute Akaike weights using the formula in (3).
Figure 8 shows the estimated Akaike weights by category and data type. We find considerable variation across categories in terms of which types of data are most important for predicting substitution patterns.
Importantly, many of these results are difficult to predict in advance. For example, while we might expect visual features to be most relevant in clothing categories, where consumers care about visual design, only two of the five clothing categories (“Activewear” and “Sleep”) assign more weight to images than to text. In contrast, for “Tops & Blouses” and “Underwear,” product titles are most predictive of substitution, whereas for “Shirts,” reviews perform best. Similarly, while we may expect descriptions and reviews of video games to be as informative as those for books, the data strongly suggests that in the category of video games for Xbox, images contain substantially more information about substitution patterns than text.
These results highlight that researchers cannot reliably predict in advance which data types will best capture substitution. Therefore, in practice, it is important to collect different data types and use model selection to identify the one that is most predictive of substitution.
@@ 5.1 Data Choice and Model Selection
The main takeaway from our results is that text and image data contain valuable information about substitution patterns, making these unstructured data useful for demand estimation. At the same time, our multi-category analysis in Section 4 shows that researchers may not be able to predict in advance which data type will best capture substitution in a given category. We therefore recommend collecting various types of unstructured data and performing model selection as in Sections 3 and 4. Algorithm 1 in Section 2 provides a practical heuristic for model selection: in our experiment, it successfully identified the specification that delivered the best counterfactual second-choice predictions.
A natural question is whether researchers should add observed attributes to our approach when they are available. In our application, adding observed attributes to the selected review-based model did not improve fit (see Section 3.3). One explanation is that unstructured data may already encode the choice-relevant attributes, rendering structured attributes redundant. Nevertheless, in other applications, researchers may still want to test whether including observed attributes helps estimate substitution patterns. This can be done by checking whether adding some subset of observed attributes to the selected model with unstructured data further reduces $AIC$. When many attributes are available, researchers can reduce their dimensionality using PCA, as in Section 4.2.
@@ 5.2 Counterfactual Analysis and its Boundaries
We interpret embeddings as capturing time-invariant product aspects (e.g., attributes, functional benefits, and visual designs). This assumption is reasonable given that in our application, product images, titles, and descriptions rarely change over time (see Appendix F). Thus, our method is well-suited for a wide range of counterfactuals that require estimating consumer responses to price changes conditional on the other product attributes, including optimal pricing, simulating merger effects on prices, evaluating corrective taxes, or estimating markups (Berry and Haile, 2024). In such contexts, researchers can evaluate counterfactual market shares and prices just as in standard demand models, assuming that embeddings—and the resulting demand functions—do not change. Moreover, as usual, researchers can quantify the welfare effects of price changes by calculating the area under the demand curve (Small and Rosen, 1981). The fact that principal components have no direct interpretation does not preclude such analysis.
Our approach can also be applied to study consumer responses to changes in product availability, for example, to quantify consumer surplus from new products or study how firms optimize assortments. In retrospective analyses, where the introduction of new products is observed in the data as in Petrin (2002), researchers can use our approach “as is” knowing that it performs well at predicting consumer responses to product removals. By contrast, evaluating a hypothetical new product that is not yet in the data poses an additional challenge: researchers need to obtain images and textual descriptions of this product from pre-launch information.
In some applications, embeddings may change in counterfactual scenarios. For example, if customer reviews mention value relative to price (e.g., “Great quality for the price!” or “Good product, but overpriced.”), then the reviews and the resulting text embeddings will shift when prices change in counterfactuals. Researchers can diagnose this issue by checking how common such reviews are in their data and, if necessary, model how price changes affect embeddings.
Another example is merger simulations: if firms redesign or reposition products after merging, this may alter embeddings and hence substitution patterns. If this is a first-order concern, researchers may want to specify a supply-side model in which firms choose both prices and non-price characteristics (Fan, 2013).
That said, our method does allow for correlation between embeddings and variables that change in counterfactuals. For instance, embeddings may be correlated with prices if they capture unobserved product quality. This does not rule out counterfactuals where prices change (e.g., due to tariffs), since such changes do not alter the product’s intrinsic quality.
@@ 5.3 Future Research Directions
Second, future research should provide a formal treatment of how to perform inference in a way that accounts for the uncertainty in the model selection step. One possibility would be to use sample-splitting as in Wasserman and Roeder (2009) and Taylor and Tibshirani (2015), randomly selecting a subset of the data for model selection and using the other for estimation and inference based on the selected model.
Third, other demand models may leverage text and image data more effectively than the random coefficients logit model. In an earlier version of this paper, we tested the pairwise combinatorial logit model of Koppelman and Wen (2000), which allows each product pair to have its own utility correlation based on distance in the embedding space. The model, however, performed significantly worse than our mixed logit, not only when using unstructured data but also with observed attributes, where we let utility correlations depend on the distance between products in the attribute space. This highlights how the functional form assumptions of the demand model can significantly impact counterfactual performance and suggests that combining unstructured data with more flexible demand models, such as that in Compiani (2022), may be a fruitful direction for future research. We note that this point applies more generally to any discrete choice model and is not specific to our method.
Finally, because our focus is on recovering substitution patterns, we largely abstract away from price endogeneity. Endogeneity is not a concern in our experiment because we randomize prices. In the Amazon application, we rely on product fixed effects to capture unobserved attributes (e.g., quality) that may correlate with prices. More generally, researchers may worry that prices correlate with unobserved demand shocks that vary across markets and over time. To address this, one could combine our approach with existing methods for handling price endogeneity using instrumental variables (Goolsbee and Petrin, 2004; Berry et al., 2004). [^28] For example, one could estimate product-market fixed effects and random coefficient variances via maximum likelihood, and then estimate a two-stage least squares regression of the estimated fixed effects on prices and principal components. The principal components of competing products could serve as excluded instruments for prices.
@@ 6 Conclusion
In this paper, we demonstrate how researchers can incorporate unstructured text and image data in demand estimation to recover substitution patterns. Our approach extracts low-dimensional features from product images and textual descriptions, integrating them into a standard mixed logit model with random coefficients. Using experimental data, we show that our approach outperforms standard attribute-based models in counterfactual predictions of second choices. We further validate our method with e-commerce data across dozens of categories and find that text and image data consistently help identify close substitutes within each category.
@@ Appendix A Text Processing Steps
Before applying our text models, listed in Section 2, we pre-process our text data as follows. Working with product titles is straightforward because each title is a short text that typically includes only 10-15 words. When assessing similarity based on product descriptions, we merge the text from all bullet points, and apply our models to the merged text. For customer reviews, we transform the text of each review into a separate vector of word occurrences or an embedding, and we average these vectors or embeddings across all reviews of a given product.
For both bag-of-words approaches, we further pre-process text data by removing stopwords and lemmatizing words. We remove stopwords using the standard dictionary of common English words in the NLTK package. We lemmatize words using the WordNet Lemmatizer from the same package, NLTK. Then, we convert each pre-processed text into a vector of word occurrences (weighted word occurrences for TF-IDF). The other two models, USE and ST, have a built-in text pre-processing step. We therefore apply these models directly to the unprocessed text data.
@@ Appendix B Book Selection Procedure
Which books we show to participants can significantly affect the performance of different demand models. Our general strategy was to select a set of books that does not bias our model comparisons in favor of any particular specification. To this end, we aimed to maximize variation in choice-relevant structured attributes as well as in text and image embeddings.
@@
To avoid arbitrary book selection within each genre, we implemented the following algorithm. We considered all possible combinations of ten books such that the three genres are roughly equally represented. Among these, we choose the combination of books that maximized the variance of text and image embeddings. [^29] Operationally, we computed the variance of image embeddings (using the VGG19 model) and the variance of text embeddings (using the USE model), then averaged the two. We performed a brute-force search over all possible sets of ten books and selected the set that maximized this average variance. The final set, shown in Figure 3, includes four mystery books, three fantasy books, and three self-help books.
@@ Appendix C Choice Survey: Sanity Checks
Because choices in our experiment were hypothetical and not incentivized, we want to verify that participants made meaningful choices. Several summary statistics suggest that participants took the choice tasks seriously. First, only 50 of the initially recruited participants (less than 1%) completed the entire study in less than one minute and thus had to be dropped from our sample. The remaining participants spent, on average, 7 minutes on the survey overall and 1.3 minutes on the choice tasks, indicating they took their time to make careful selections and did not mindlessly click through the survey (Figure A1).
Second, in the choice tasks, participants were disproportionately more likely to select books of the genre they reported to be their favorite in a questionnaire before the choice tasks (Figure A2), suggesting they considered the book attributes when making their choices.
Finally, participants’ choices were consistent across the two choice tasks (Figure A3). For example, over 60% of participants who selected a mystery book in the first task chose another mystery book in the second choice task. This observation suggests participants considered their genre preferences when making both choices.
@@ Appendix D Computing Second-Choice Diversion Ratios
where $y_{i}^{(1)}$ and $y_{i}^{(2)}$ are participant $i$’s first and second choices. By contrast, to compute diversions $\hat{s}_{j\rightarrow k}$ predicted by a given demand model, we use the identity
@@
Appendix E Attributes Collected for Electronics Products
For each of the four electronics categories described in Section 4.1, we collect standard attributes from Amazon product detail pages. Specifically, we include all specifications listed in the product descriptions and, when available, those in the “Technical Details”’ section. Below is the complete list of all attributes for each category:
@@ Appendix E Attributes Collected for Electronics Products
Appendix F Changes in Text and Images Over Time
As discussed in Section 2, we treat text and image embeddings, as well as the principal components extracted from them, as being fixed over time for a given product. To validate this assumption, we examine whether text and images change over time.
Since we lack data on these changes for 2019-2020, we construct a separate sample by repeatedly collecting unstructured data from Amazon’s product detail pages daily from January 23 to March 4, 2025. To keep data collection manageable, we do not gather customer reviews and select a subset of 11 out of 40 categories, ensuring they cover all of Amazon’s departments (e.g., “Clothing,” “Food,” “Electronics”) represented in the full dataset. [^30] Recall that we average over embeddings extracted from the 100 most recent reviews. Even though customer reviews accumulate over time as consumers continue to write them, the average embeddings extracted from these reviews may remain approximately constant if consumers consistently discuss the same attributes. While we do not have review data to verify this claim, future research should examine whether review embeddings are stable in online markets. The selected categories are: “Shirts,” “Coffee,” “Aromatherapy,” “Mattresses,” “Markers,” “Pet Litters,” “Nintendo Games,” “Tablets,” “Memory Cards,” “Monitors,” and “Earbuds.” In each category, we collect data for the products used in our estimation in Section 4 that were not discontinued, totaling 136 products.
We find that product images do not change over time. Titles change for only six products (4%), mostly by adding or removing specific attributes or functional benefits. Similarly, descriptions change for just 21 products (15%). Most changes do not affect which product features are revealed, but they alter which ones are immediately visible versus being revealed only after clicking on “See more product details.” Thus, we conclude that changes in text and images are not a significant concern for the products in our empirical application.

### v4

@@ 4.3 Relevance of Different Data Types
Lastly, we analyze which types of unstructured data yield the largest fit improvements. It would be misleading to report only the best-performing model in each product category, as multiple text or image models may achieve similar $AIC$ improvements over plain logit. We therefore follow the statistical literature on model selection and report Akaike weights (Burnham and Anderson, 2004). Formally, let $d$ denote data type (images, titles, descriptions, or reviews), and let $AIC_{d}$ denote the lowest $AIC$ among all specifications that use data type dd. [^21] The summation in the denominator is over all four data types. For each data type $d$, we compute the Akaike weight $w_{d}$ as
$w_{d}=\frac{\exp(-{}_{d}/2)}{\sumop\displaylimits_{k}\exp(-{}_{k}/2)}$ (3)
where $_{d}=AIC_{d}-AIC_{min}$ and $AIC_{min}=\min_{d}AIC_{d}$ is the lowest $AIC$ across all four data types. We interpret $w_{d}$ heuristically as the posterior probability that the mixed logit model based on data type $d$ is the best model given the data. [^22] More precisely, in large samples, $w_{d}$ reflects the probability that this class of models is, in fact, the best model for the data in the sense of Kullback-Leibler information (Burnham and Anderson, 2004, p.272).
Figure 7: Akaike weights of different types of unstructured data across 40 product categories. These weights reflect the relative performance of data types in capturing substitution patterns from choice data. We compute Akaike weights using the formula in (3).
Figure 7 shows the estimated Akaike weights by category and data type, revealing substantial variation across categories. Many results are difficult to anticipate a priori. For example, although one might expect visual features to be most relevant for clothing, only two of the five clothing categories (“Activewear” and “Sleep”) place more weight on images than text. In “Tops & Blouses” and “Underwear,” product titles are most predictive of substitution, while in “Shirts,” reviews perform best. Similarly, we may expect descriptions and reviews of video games to be most informative, analogous to books. In practice, we find that images contain substantially more information about substitution in the “Video Games Xbox” category.
These results highlight that researchers cannot reliably predict which data types best capture substitution. In practice, it is therefore important to collect multiple data types and use model selection to choose the one that best captures substitution.
@@ 5.1 Data Choice and Model Selection
The main takeaway from our results is that text and image data contain valuable information about substitution patterns, making these unstructured data useful for demand estimation. At the same time, our multi-category analysis in Section 4 shows that researchers may not be able to predict in advance which data type will best capture substitution in a given category. We therefore recommend collecting various types of unstructured data and performing model selection as in Sections 3 and 4. Algorithm 1 in Section 2 provides a practical heuristic for model selection: in our experiment, it successfully identifies the specification that produces the best counterfactual second-choice predictions.
A natural question is whether researchers should add observed attributes to our approach when they are available. In our application, adding observed attributes to the selected review-based model does not improve fit (see Section 3.3). One explanation is that unstructured data may already encode the choice-relevant attributes, rendering structured attributes redundant. Nevertheless, in other applications, researchers may still want to test whether including observed attributes helps estimate substitution patterns. For example, researchers can expand the set of candidate variables in Algorithm 1 to include the observed attributes themselves or, when many attributes are available, their principal components.
@@ 5.2 Counterfactual Analysis and its Boundaries
Our approach is well suited for applications that assess consumer responses to price changes while holding embeddings fixed, including optimal pricing, merger simulations, corrective taxes, and markup estimation (Berry and Haile, 2024). In these applications, embeddings may be correlated with prices in the observed data, but they, and hence the demand system, should be held fixed in counterfactual simulations. Under this assumption, researchers can compute counterfactual market shares, prices, and welfare using standard tools from demand analysis. For instance, consumer welfare can be computed as the area under the demand curve (Small and Rosen, 1981). The fact that embeddings have no direct interpretation does not preclude these analyses.
Our approach can also predict responses to changes in product availability, for example, to study firms’ assortment choices or consumer surplus from new products (Petrin, 2002). If new products launches are observed in the data, researchers can apply our method directly, since our results show it performs well at predicting consumer responses to product removals. On the other hand, when evaluating a hypothetical new product, researchers must construct its embeddings from pre-launch information, such as images and descriptions from marketing materials or early retailer listings.
By contrast, our approach is less suitable in applications where embeddings may change in counterfactuals. For example, if customer reviews mention value relative to price (e.g., “Great quality for the price!” or “Good product, but overpriced.”), embeddings constructed from them may shift when prices change. Researchers can diagnose this issue by checking how frequently price-related language appears in reviews and, if necessary, modeling how price changes affect embeddings. Another example is simulations of mergers that may induce firms to redesign products, which would alter embeddings and hence substitution patterns. Studying such mergers is not possible with our approach unless researchers impose additional assumptions on firm conduct and on how embeddings map into attributes that firms can control.
@@ 5.3 Future Research Directions
Second, our method incorporates embeddings into the demand model as if they were the true attributes driving substitution, when in practice embeddings may be only imperfect proxies of these attributes. Our results show that, despite this limitation, embeddings are sufficiently informative to improve counterfactual predictions. Nevertheless, future research should explore whether explicitly accounting for error in these proxies can further improve performance. For an early contribution in this area, see Christensen and Compiani (2026).
Third, other demand models may leverage unstructured data more effectively. In an earlier version of this paper, we tested the pairwise combinatorial logit model of Koppelman and Wen (2000), which allows pairwise utility correlations to depend on distances in the embedding space. That model performed worse than mixed logit, not only when using unstructured data but also when using distances in observed attributes, suggesting that functional form assumptions can significantly impact counterfactual performance. A fruitful direction for future research might be to combine unstructured data with more flexible demand models, such as that in Compiani (2022).
Finally, because our focus is on recovering substitution patterns, we largely abstract away from price endogeneity. Endogeneity is not a concern in our experiment because we randomize prices. In the e-commerce application, we rely on product fixed effects to capture unobserved attributes (e.g., quality) that may correlate with prices. More generally, prices may correlate with unobserved demand shocks that vary across markets and over time. To address this, one could combine our approach with existing methods for handling price endogeneity using instrumental variables (Goolsbee and Petrin, 2004; Berry et al., 2004). [^24] For example, one could estimate product-market fixed effects and random coefficient variances via maximum likelihood, and then estimate a two-stage least squares regression of the estimated fixed effects on prices and principal components. The principal components of competing products could serve as excluded instruments for prices.
@@ 6 Conclusion
In this paper, we show how researchers can incorporate unstructured text and image data in demand estimation to recover substitution patterns. Our approach extracts low-dimensional features from product images and textual descriptions, integrating them into a standard mixed logit model with random coefficients. Using experimental data, we show that our approach outperforms standard attribute-based models in counterfactual predictions of second choices. We further validate our method with e-commerce data across dozens of categories and find that text and image data consistently help identify close substitutes within each category.
@@ Appendix A Text Processing Steps
Before computing embeddings, we pre-process text as follows. We keep product titles unchanged because they typically consist of only 10-15 words. For product descriptions, we merge the text from all bullet points and apply the embedding models to the merged text. For customer reviews, we transform the text of each review into a vector of word occurrences or an embedding, and we average these vectors or embeddings across all reviews of a given product.
For both bag-of-words models, we further pre-process text data by removing stopwords and lemmatizing words. Stopwords are removed using the standard dictionary of common English words in the NLTK package. We lemmatize words using the WordNet Lemmatizer from the same package, NLTK. We then convert each processed text into a vector of word occurrences (or weighted word occurrences for TF-IDF). By contrast, the USE and ST models have built-in text pre-processing, so we apply them directly to the raw text.
@@ Appendix B Book Selection Procedure
The set of books shown to participants can significantly affect the performance of demand models. Our goal was to select books in a way that does not favor any particular specification. To this end, we maximized variation in both observed attributes as well as in text- and image-based embeddings.
@@
To avoid arbitrary book selection within each genre, we implemented the following algorithm. We considered all possible combinations of ten books such that the three genres are roughly equally represented. Among these, we choose the combination of books that maximized the variance of text and image embeddings. [^25] We computed the variance of image embeddings (using the VGG19 model) and the variance of text embeddings (using the USE model), then averaged the two. We performed a brute-force search over all possible sets of ten books and selected the set that maximized this average variance. The final set, shown in Figure 3, includes four mystery books, three fantasy books, and three self-help books.
@@ Appendix C Choice Survey: Sanity Checks
Because choices in our experiment were not incentivized, it is crucial to verify that participants made meaningful choices. Several summary statistics suggest that participants took the choice tasks seriously. First, only 50 of the initially recruited participants (less than 1%) completed the entire study in under one minute and thus were excluded from the sample. The remaining participants spent, on average, 7 minutes on the survey overall and 1.3 minutes on the choice tasks, indicating they took their time to make careful selections and did not mindlessly click through the survey (Figure A2).
Second, participants were disproportionately more likely to select books of the genre they reported to be their favorite in a questionnaire before the choice tasks (Figure A3). Further, their choices were consistent across the two choice tasks (Figure A4). For example, over 60% of participants who selected a mystery book in the first task chose another mystery book in the second choice task. These observations suggest that participants considered their genre preferences when making choices.
@@ Appendix D Computing Second-Choice Diversion Ratios
where $y_{i}^{(1)}$ and $y_{i}^{(2)}$ are participant $i$’s first and second choices. By contrast, to compute diversions $\hat{s}_{j\rightarrow k}$ predicted by a given demand model, we use the following identity from Conlon and Mortimer (2021):
@@
Appendix E Implications for Pricing: Merger Simulations
To illustrate how estimated substitution patterns can influence counterfactuals of interest, we conduct a stylized simulation of horizontal mergers—a natural application of our approach. Antitrust agencies routinely use demand models to assess whether a hypothetical merger would lead to a significant price increase (Federal Trade Commission, 2022). If the merging firms’ products are close substitutes, the merger creates strong upward pricing pressure, as the firm can “recapture” some consumers after raising prices. Therefore, predicting how a merged firm would set prices requires accurate estimates of substitution patterns.
Figure A1: Results of hypothetical merger simulations. For each simulated merger of Dopamine Detox with one other book (horizontal axis), the figure shows the predicted average price increase among the two merging books. The horizontal dashed line represents a hypothetical policy where the decision-maker challenges all mergers expected to raise prices by at least 5%.
We perform simulations using our demand estimates from Section 3. For each pair of books, we compute their prices under two scenarios: (a) when the books are owned by separate publishers competing in a Bertrand-Nash equilibrium, and (b) when both books are owned by the same publisher setting prices by solving a joint first-order condition. In both cases, publishers take the prices of the other eight books as given, fixed at \$5, and have zero marginal costs. [^26] We fix the prices of the other books at \$5, the average value in the experimental dataset. We do not optimize prices of other books because the model does not include an outside option. Consumer choices remain unchanged if all prices increase by the same amount, thus leading to multiple equilibria.
We recognize that restricting the choice set to ten books and excluding an outside option makes the setting somewhat artificial. Still, we view it as a useful way to evaluate how estimated substitution patterns translate into counterfactuals of interest.
Figure A1 illustrates the predicted price increase resulting from the joint ownership of Dopamine Detox with each of the other books. We select Dopamine Detox as an example where our approach outperforms alternative models in capturing substitution patterns (see Table 2). For each simulated merger, Figure A1 reports the average price increase across the two merging books.
As discussed in Section 3, the plain logit model fails to capture strong within-genre substitution. Consequently, it incorrectly identifies as the closest substitutes for Dopamine Detox the three mystery books, rather than other self-help ones. This misclassification affects the predicted price changes in merger simulations. Specifically, the plain logit model overstates the price increase when Dopamine Detox is merged with the mystery book Please Tell Me or The Housemaid, while understating the price increase when merged with its true closest substitutes, The Art of Letting Go or Don’t Believe Everything You Think. These discrepancies are substantial: for within-genre mergers, the plain logit predicts modest price increases of 3% and 3.7%, whereas our review-based model estimates them to be 5.6% and 8.9%—approximately 2 to 2.5 times higher.
Appendix F Attributes Collected for Electronics Products
For each of the four electronics categories described in Section 4.1, we collect standard attributes from Amazon product detail pages. Specifically, we include all specifications listed in the product descriptions and, when available, those in the “Technical Details” section of the product detail page. Below is the complete list of all attributes for each category:
@@ Appendix F Attributes Collected for Electronics Products
Appendix G Changes in Text and Images Over Time
Throughout the paper, we treat text and image embeddings, as well as the principal components extracted from them, as being fixed over time for a given product. To validate this assumption, we examine whether text and images change over time.
Since we lack data on these changes for 2019-2020, we construct a separate sample by repeatedly collecting unstructured data from Amazon’s product detail pages daily from January 23 to March 4, 2025. To keep data collection manageable, we do not gather customer reviews and select a subset of 11 out of 40 categories, ensuring they cover all of Amazon’s departments (e.g., “Clothing,” “Food,” “Electronics”) represented in the full dataset. [^27] Recall that we average over embeddings extracted from the 100 most recent reviews. Even though customer reviews accumulate over time as consumers continue to write them, the average embeddings extracted from these reviews may remain approximately constant if consumers consistently discuss the same attributes. While we do not have review data to verify this claim, future research should examine whether review embeddings are stable in online markets. The selected categories are: “Shirts,” “Coffee,” “Aromatherapy,” “Mattresses,” “Markers,” “Pet Litters,” “Nintendo Games,” “Tablets,” “Memory Cards,” “Monitors,” and “Earbuds.” In each category, we collect data for the products used in estimation in Section 4 that were not discontinued, totaling 136 products.
We find that product images do not change over time. Titles change for only six products (4%), mostly by adding or removing specific attributes or functional benefits. Similarly, descriptions change for just 21 products (15%). Most changes do not affect which product features are revealed, but they alter which ones are immediately visible versus being revealed only after clicking on “See more product details.” Thus, we conclude that changes in text and images are not a significant concern in our empirical application.